INFRARED SPECTROSCOPY OF COLD, HYDRATED ALKALINE EARTH-HYDROXIDE CLUSTERS

\underline{CHRISTOPHER JOHNSON}, CHRISTOPHER LEAVITT, JOSEPH FOURNIER, MARK JOHN-SON, Department of Chemistry, Yale University, New Haven, CT 06520.

Hydrated, singly-charged [MOH]^{+} (M=Mg,Ca) clusters have been investigated using mass-selective cryogenic infrared spectroscopy. Spectra of [MOH]^{+}· (H_{2}O)_{n} reveal broad features beginning at n = 3 that approach 1000 cm^{−1} in width at n = 5, indicative of large amplitude motion in the water network despite ion temperatures below 40 K. Comparison to calculated structures and spectra help to elucidate the structural and charge-separation dynamics occurring in these clusters.